OpenStudy (anonymous):

A company sells boxes of small blue pom-poms (b) for $42 and boxes of large orange pom-poms (p) for $52. They make batches of small blue pom-poms that fill 8 boxes and batches of large orange pom-poms that fill 4 boxes. The company only has 40 boxes. Which system of equations helps the company plan to make $350?

OpenStudy (anonymous):

a. b = 40 - p 8b + 4p = 40 b. b = 40 - p 42b + 52p = 350 c. 8b = 40 - 4p 42b - 52p = 350 d. 8b = 40 - 4p 42b + 52p = 350

OpenStudy (anonymous):

b= 42 p = 52 8b + 4p = 40 42b + 52p = 350 This is what I had but I don't feel confident with it. Just typing it as an opinion as of now

OpenStudy (anonymous):

It looks like d but I didn't like the wording in the problem If they only sell the boxes by 8 small and 4 large then yes

OpenStudy (anonymous):

I think your right.

OpenStudy (anonymous):

thanks for trying